# Generative Enzyme Design Guided by Functionally Important Sites and Small-Molecule Substrates

### v3

@@ 3.2 NAEL Global Attention Sub-Layer
where $\boldsymbol{h}_{i}^{l}$ is the $i^{th}$ residue input representation at $l^{\mathrm{th}}$ layer and $\boldsymbol{H}_{e}^{l}=[\boldsymbol{h}_{1}^{l},\boldsymbol{h}_{2}^{l},...,\boldsymbol{h}_{N}^{l}]^{T}$. The input residue embeddings for the first layer are either taken from an embedding lookup table for functionally important sites, or initialized with a special [mask] token embedding for other residues. The residue embedding will then be enhanced by the enzyme family embedding Emb($c$) corresponding to the desired enzyme family $c$:
$\small\begin{split}\boldsymbol{h}_{i}^{0}&=\hat{\boldsymbol{h}}_{i}^{0}+\text{Emb}(c)\\ \hat{\boldsymbol{h}}_{i}^{0}&=\left\{\begin{aligned} W_{\mathcal{A}}^{T}\boldsymbol{s}_{i}&,&{\text{$i\in\mathcal{M}$,}}\\ \text{Emb([mask]})&,&{\text{otherwise}}\end{aligned}\right.\end{split}$ (6)
@@ 3.3 NAEL Neighborhood Equivariant Sub-Layer
$\footnotesize\begin{split}\boldsymbol{m}_{ik}^{l+0.5}&=\mathrm{SiLU}(\mathrm{FFN}([(\boldsymbol{h}_{i}^{l+0.5};\boldsymbol{h}_{k}^{l+0.5};||\boldsymbol{x}_{i}^{l}-\boldsymbol{x}^{l}_{k}||_{2}]))\\ w_{ik}^{l+0.5}&=\frac{\exp(W_{a}^{l}\boldsymbol{m}_{ik}^{l+0.5}+b_{a}^{l})}{\sum_{k^{\prime}\in\mathrm{Neighbor(i)}}\exp(W_{a}^{l}\boldsymbol{m}_{ik^{\prime}}^{l+0.5}+b_{a}^{l})}\\ \boldsymbol{m}_{ik}^{l+1}&=w_{ik}^{l+0.5}*\boldsymbol{m}_{ik}^{l+0.5}\\ \end{split}$ (7)
where FFN is a two-layer fully connected feed-forward network with SiLU activation function after its first layer. [;] is concatenation operator and $||\boldsymbol{x}_{i}^{l}-\boldsymbol{x}^{l}_{k}||_{2}$ is the Euclidean distance between $i^{th}$ and $k^{th}$ residue coordinates at $l^{th}$ layer. $W_{*}$ and $b_{*}$ are trainable parameters.
@@
The input residue $C_{\alpha}$ coordinates are either the given coordinates for functionally important sites, or randomly initialized as 3D points on the spherical surface centered at its left residue, considering the Euclidean distances between neighboring $C_{\alpha}$ pairs are almost the same (around $r=3.75$Å):
$\small\boldsymbol{x}_{i}^{0}=\left\{\begin{aligned} \boldsymbol{x}_{i},&\text{$i\in\mathcal{M}$},\\ \boldsymbol{x}_{i-1}^{0}+r\cdot[\sin{\omega_{1}}\cos{\omega_{2}},\sin{\omega_{1}}\sin{\omega_{2}},\cos{\omega_{1}}]^{T},&\text{$i\notin\mathcal{M}$}\end{aligned}\right.$ (9)
@@
$\footnotesize\begin{split}\boldsymbol{c}_{i}^{l+1}&=\sum\nolimits_{k\in\mathrm{Neighbor(i)}}\boldsymbol{m}_{ik}^{l+1}\\ \boldsymbol{h}_{i}^{l+1}&=\boldsymbol{h}_{i}^{l+0.5}+\sigma(\mathrm{FFN}(\boldsymbol{c}_{i}^{l+1}))\odot\boldsymbol{c}_{i}^{l+1}\end{split}$ (10)
@@ 3.4 Substrate Representation Module
$\small\begin{split}\boldsymbol{m}_{jk}^{l+1}&=\phi_{m}(\boldsymbol{h}_{j}^{l},\boldsymbol{h}_{k}^{l},||\boldsymbol{x}_{j}^{l}-\boldsymbol{x}^{l}_{k}||_{2})\\ \boldsymbol{h}_{j}^{l+1}&=\boldsymbol{h}_{j}^{l}+\phi_{n}(\boldsymbol{h}_{j}^{l},\sum\nolimits_{k\in\mathrm{Neighbor(j)}}\boldsymbol{m}_{jk}^{l+1})\end{split}$ (12)
where $\phi_{m}$ and $\phi_{n}$ respectively denotes neighborhood message update and neighborhood node feature update. Following Guan et al. (2022), we initialize the node feature as five chemical features $\boldsymbol{\hat{h}}_{j}^{0}\in\mathbb{R}^{5}$ and $\boldsymbol{h}_{j}^{0}=W_{s}\cdot\boldsymbol{\hat{h}}_{j}^{0}$ where $W_{s}\in\mathbb{R}^{d\times 5}$ is a mapping matrix and $d$ is the representation dimensionality. Since we have the real 3D structure of the substrate, we do not need to update its atom coordinates.
@@ 3.5 Functionally Important Enzyme Site Discovery
Fourth-Level Category 3157 428 323
@@ 4 Experiments
Table 2: ESP scores ($\uparrow$) of generated enzymes for 30 testing BRENDA EC categories. ESP evaluates the enzyme-substrate interaction function. Each number is the average ESP score of 50 testing cases for each category. IF denotes the inverse folding model ProteinMPNN. A ESP score of 0.6 or higher indicates a positive substrate binding. The last column (Avg) is the average across 30 categories. Notice that our EnzyGen outperforms all previous methods in ESP by a big margin.
Protein Family 1.1.1 1.11.1 1.14.13 1.14.14 1.2.1 2.1.1 2.3.1 2.4.1 2.4.2 2.5.1 2.6.1 2.7.1 2.7.10 2.7.11 2.7.4
@@
4.1 EnzyBench Construction
We aim to evaluate the effectiveness of our EnzyGen across a diverse range of enzymes. To achieve this, we gather all enzymes from BRENDA (Schomburg et al., 2002), yielding a total of $101,974$ PDB entries. These entries are classified into $3,157$ fourth-level enzyme categories based on the EC Tree (Appendix Figure 5). It is noteworthy that there exist a total of $8,422$ fourth-level categories; however, our focus lies specifically on those with experimentally confirmed structures. This focus results in our dataset comprising $3,157$ fourth-level categories. We then conduct MSAs for each fourth-level category to identify functionally important sites, using an MSA threshold ($\tau$) of 30%. Given the relatively small size of fourth-level categories, we aggregate data belonging to the same third-level category in the EC Tree, resulting in $256$ third-level categories. Then we select 30 third-level categories for validation and testing, respectively including 428 and 323 fourth-level categories. For each of the 30 third-level categories, we we randomly split $100$ PDB entries with 50 for validation and 50 for testing, while the remaining entries are utilized for training. To prevent data leakage, we cluster the PDB entries with a sequence identity threshold of $50\%$. We ensure that no PDB entries in the validation or test sets belong to the same cluster as those in the training set. We further collect $23,711$ experimentally confirmed $<$enzyme, substrate$>$ pairs from Kroll et al. (2023), ensuring enzymes without a corresponding substrate have a randomly sampled negative one during the training process. We also make sure that all entries in the test set have a corresponding substrate to ensure function evaluation. Detailed data statistics are provided in Table 1 and Appendix C.
4.2 Experimental Setup
Implementation Details. We use 33 NAELs to model enzymes and 3 neighborhood equivariant layers for substrate representation module with node feature dimensionality set to 1280. As suggested by Ying et al. (2018), the number of graph layers typically ranges from 2 to 6. To prevent over-fitting and facilitate efficiency, we utilize 3 neighborhood equivariant sub-layers for enzyme modeling, comprising a modified NAEL with every 11 global attention sub-layers followed by one neighborhood equivariant sub-layer. The parameters of the global attention sub-layer are initialized with the released 650M ESM-2 (Lin et al., 2022b) parameters. The total parameter count is 714M. The hyperparameters $\lambda/2$ and $K$ are set to $1.0$ and $30$, respectively. The model undergoes training for $1,000,000$ steps using 8 NVIDIA RTX A6000600060006000 GPU cards, among which only the sequence generation and position prediction losses are adopted for the first $200,000$ steps. The batch size and learning rate are set to $8192$ tokens and $3$e-$4$ respectively. Sequences are decoded using the greedy decoding strategy. We incorporate four-level enzyme tags in the EC tree. For example, an enzyme in the category 1.1.1.1 should have four tags: 1, 1.1, 1.1.1, and $1.$
@@ 4.3 Main Results
EnzyGen excels in designing enzymes that bind their respective substrates with high affinities. Table 2 and 3 highlight the ability of EnzyGen in achieving the highest average ESP score and substrate binding affinity across 323 families. Notably, EnzyGen obtains an average ESP score of $0.65$, surpassing the suggested enzyme-substrate interaction threshold of 0.6 from the ESP evaluator developer. In comparison, PROTSEED achieves 0.59, RFDiffusion+ProteinMPNN achieves 0.53, and ESM2+EGNN achieves 0.61 of average ESP score. EnzyGen’s ESP score is significantly higher than all previous baselines. This is further verified by the binding affinity scores. The enzymes generated by our EnzyGen exhibit stronger binding affinity to their corresponding substrates than all other methods across all 30 third-level testing EC categories. EnzyGen gains an average improvement of $-0.87$ in binding affinity over the previous best method (RFDiffusion+ProteinMPNN), which is significant. These results demonstrate that enzymes designed by our EnzyGen are able to bind their corresponding substrates with high affinities.
5 Analysis: Diving Deep into EnzyGen
(a) Zero-shot generalization
(b) Fourth-level category clustering
(c) Complex of 1KAG-ATP(-4)
(d) Complex of 5L2P-paraoxon
Figure 4: (a) ESP scores of designed enzymes from new fourth-level classes or with new substrates. Dash line denotes median. (b) Fourth-level category embedding clustering. Case Study: (c) Complex of designed 1KAG (2.7.1.71, catalyzing the specific phosphorylation of the 3-hydroxyl group of shikimic acid) and substrate ATP(-4), with pLDDT=90.39, Uniprot blastp recovery rate = 58.5%, (d) Complex of designed 5L2P (3.1.1.2, hydrolyzing various p-nitrophenyl phosphates, aromatic esters and p-nitrophenyl fatty acids) and substrate paraoxon, with pLDDT=89.44, Uniprot blastp recovery rate = 49.4%. Both cases show polar contacts (hydrogen bonds) depicted in purple.
Protein Family 1.1.1 1.11.1 1.14.13 1.14.14 1.2.1 2.1.1 2.3.1 2.4.1 2.4.2 2.5.1 2.6.1 2.7.1 2.7.10 2.7.11 2.7.4
@@ 5.6 Designing Novel Enzymes with EnzyGen
EnzyGen is able to design novel enzymes that bind to specific substrates. Figure 4 (c) and (d) showcase two designed enzymes based on PDB 1KAG (catalyzing the specific phosphorylation of the 3-hydroxyl group of shikimic acid) and 5L2P (hydrolyzing various p-nitrophenyl phosphates, aromatic esters and p-nitrophenyl fatty acids) respectively. Both of the two designed enzymes achieve pLDDT approaching or exceeding 90 and substrate binding affinities below -10. Additionally, the Gnina docked complexes show polar contacts (hydrogen bonds) between the enzymes and substrates. This affirms the enzyme-substrate interaction function. We carry out blastp search of designed enzymes in Uniprot, yielding an amino acid identity of 58.5% and 49.4%, respectively, to the most similar enzyme. Together, these results show that our model is capable of designing novel enzymes with high substrate binding affinities. Additional cases are detailed in Appendix D.3, each exhibiting a pLDDT higher than 80, a binding affinity below -10, and a Uniprot blastp identity rate lower than 65%. Enzyme sequences for all analyzed cases are also available in Appendix Table 19.
@@ A.1 Proof: NAEL Is SE(3) Equivariant
For message update, the distance between two residues is invariant as $d_{ij}^{l}=||R\boldsymbol{x}_{i}^{l}+\boldsymbol{t}-(R\boldsymbol{x}^{l}_{j}+\boldsymbol{t})||^{2}=(\boldsymbol{x}_{i}^{l}-\boldsymbol{x}^{l}_{j})^{T}R^{T}R(\boldsymbol{x}_{i}^{l}-\boldsymbol{x}^{l}_{j})=||\boldsymbol{x}_{i}^{l}-\boldsymbol{x}^{l}_{j}||^{2}$. $\boldsymbol{h}^{l}_{i}$ is the embedding of residue or [mask] token, and thus it is always invariant. Therefore, $\boldsymbol{h}^{l+0.5}_{i}$ is also invariant. Since $\boldsymbol{h}^{l+0.5}_{i}$, $\boldsymbol{h}^{l+0.5}_{j}$ and $d_{ij}^{l}$ are all invariant, and thus $\boldsymbol{m}_{ij}^{l+1}$ is also invariant to translation $\boldsymbol{t}$ and rotation $R$ on $\boldsymbol{x}$.
@@
We have proved that when $l=0$, $\boldsymbol{h}^{1}_{i}$ is invariant to rotation and translation on $\boldsymbol{x}$. Taking $\boldsymbol{H}^{1}$ as the second layer input and following the above process, we can prove $\boldsymbol{h}^{2}_{i}$ is also invariant. Repeating this process from $l=1$ to $L-1$, we can get the same conclusion.
Combining the above two scenarios together, we have $\boldsymbol{H}^{l+1},R\boldsymbol{x}^{l+1}+\boldsymbol{t}=\mathrm{NAEL}(\boldsymbol{H}^{l},R\boldsymbol{x}^{l}+\boldsymbol{t})$ for $l=0$ to $L-1$. Therefore, our proposed NAEL is $SE(3)$-equivariant.

### v4

@@ 3.2 NAEL Global Attention Sub-Layer
where $\boldsymbol{h}_{i}^{l}$ is the $i^{th}$ residue input representation at $l^{\mathrm{th}}$ layer and $\boldsymbol{H}_{e}^{l}=[\boldsymbol{h}_{1}^{l},\boldsymbol{h}_{2}^{l},...,\boldsymbol{h}_{N}^{l}]^{T}$. The input residue embeddings for the first layer are either taken from an embedding lookup table for functionally important sites, or initialized with a special [mask] token embedding for other residues. The residue embedding will then be enhanced by the enzyme family embedding Emb($c$) corresponding to the desired enzyme family $c=\{c_{1},c_{2},c_{3},c_{4}\}$:
[…]
@@ 3.3 NAEL Neighborhood Equivariant Sub-Layer
[…]
where FFN is a two-layer fully connected feed-forward network with SiLU activation function after its first layer. [;] is concatenation operator and $\boldsymbol{x}_{i}^{l}-\boldsymbol{x}^{l}_{k}||_{2}$ is the Euclidean distance between $i^{th}$ and $k^{th}$ residue coordinates at $l^{th}$ layer. $W_{*}$ and $b_{*}$ are trainable parameters.
@@
The input residue $C_{\alpha}$ coordinates are either the given coordinates for functionally important sites, or randomly initialized as 3D points on the spherical surface centered at its preceding residue, considering the Euclidean distances between neighboring $C_{\alpha}$ pairs are almost the same (around r=3.75r=3.75Å):
$\small\boldsymbol{x}_{i}^{0}=\left\{\begin{aligned} &\boldsymbol{x}_{i}&\text{$i\in\mathcal{M}$}\\ &\boldsymbol{x}_{i-1}^{0}+r\cdot[\sin{\omega_{1}}\cos{\omega_{2}},\sin{\omega_{1}}\sin{\omega_{2}},\cos{\omega_{1}}]^{T}&\text{$i\notin\mathcal{M}$}\end{aligned}\right$.
@@
$\footnotesize\begin{split}\boldsymbol{g}_{i}^{l+1}&=\sum\nolimits_{k\in\mathrm{Neighbor(i)}}\boldsymbol{m}_{ik}^{l+1}\\ \boldsymbol{h}_{i}^{l+1}&=\boldsymbol{h}_{i}^{l+0.5}+\sigma(\mathrm{FFN}(\boldsymbol{g}_{i}^{l+1}))\odot\boldsymbol{g}_{i}^{l+1}\end{split}$ (9)
@@ 3.4 Substrate Representation Module
[…]
where $\phi_{m}$ and $\phi_{n}$ respectively denotes neighborhood message update Eq.(8) and neighborhood node feature update Eq.(9) with separate parameters. Following Guan et al. (2022), we initialize the node feature as five chemical features $\boldsymbol{\hat{h}}_{j}^{0}\in\mathbb{R}^{5}$ and $\boldsymbol{h}_{j}^{0}=W_{s}\cdot\boldsymbol{\hat{h}}_{j}^{0}$ where $W_{s}\in\mathbb{R}^{d\times 5}$ is a mapping matrix and $d$ is the representation dimensionality. Since we have the real 3D structure of the substrate, we do not need to update its atom coordinates.
@@ 3.5 Functionally Important Enzyme Site Discovery
Fourth-Level Category 3135 428 323
@@ 4 Experiments
EnzyGen-1.5 0.80 0.75 0.56 0.86 0.45 0.25 0.68 0.58 0.20 0.67 0.77 0.40 0.45 0.93 0.55 0.67
Table 2: Enzyme-substrate interaction (ESP) scores ($\uparrow$) of generated enzymes for 30 testing BRENDA EC categories averaged over 50 cases in each category. IF denotes the inverse folding model ProteinMPNN. A ESP score of 0.6 or higher indicates a positive substrate binding. The last column (Avg) is the average across 30 categories. Notice that our EnzyGen outperforms all previous methods in ESP by a big margin.
Enzyme Family 1.1.1 $1.1$ 1.14.13 1.14.14 1.2.1 2.1.1 2.3.1 2.4.1 2.4.2 2.5.1 2.6.1 2.7.1 2.7.10 2.7.11 2.7.4
@@
4.1 EnzyBench Construction
We aim to evaluate the effectiveness of our EnzyGen across a diverse range of enzymes. To achieve this, we gather all enzymes from BRENDA (Schomburg et al., 2002), yielding a total of $101,974$ PDB entries. These entries are classified into $3,157$ fourth-level enzyme categories based on the EC Tree (Appendix Figure 7). It is noteworthy that there exist a total of $8,422$ fourth-level categories; however, our focus lies specifically on those with experimentally confirmed structures. This focus results in our dataset comprising $3,157$ fourth-level categories. We then conduct MSAs for each fourth-level category to identify functionally important sites, using an MSA threshold ($\tau$) of 30%. Given the relatively small size of fourth-level categories, we aggregate data belonging to the same third-level category in the EC Tree, resulting in $256$ third-level categories. Then we select 30 third-level categories for validation and testing, respectively including 428 and 323 fourth-level categories. For each of the 30 third-level categories, we we randomly split $100$ PDB entries with 50 for validation and 50 for testing, while the remaining entries are utilized for training. To prevent data leakage, we cluster the PDB entries with a sequence identity threshold of $50\%$. We ensure that no PDB entries in the validation or test sets belong to the same cluster as those in the training set. We further collect $23,711$ experimentally confirmed <<enzyme, substrate>> pairs from Kroll et al. (2023), ensuring enzymes without a corresponding substrate have a randomly sampled negative one during the training process. We also make sure that all entries in the test set have a corresponding substrate to ensure function evaluation. Detailed data statistics are provided in Table 1 and Appendix C.
4.2 Experimental Setup
Implementation Details. We use 33 NAELs to model enzymes and 3 neighborhood equivariant layers for substrate representation module. The node feature dimensionality is set to 1280. As suggested by Ying et al. (2018), the number of graph layers typically ranges from 2 to 6. To prevent over-fitting and facilitate efficiency, we utilize 3 neighborhood equivariant sub-layers for enzyme modeling. The resulting EnzyGen consists of a modified NAEL with every 11 global attention sub-layers followed by one neighborhood equivariant sub-layer. The parameters of the global attention sub-layer are initialized with the released 650M ESM-2 (Lin et al., 2022b) parameters. The total number of parameter is 714 million. The variance and neighbor hyperparameters $\lambda/2$ and $K$ are set to $1.0$ and $30$. The model undergoes training for $1,000,000$ steps using 8 NVIDIA RTX A60006000 GPUs. The model is trained with only the sequence generation and position prediction losses for the first $200,000$ steps, and then continues training on the sequence generation, position prediction and enzyme-substrate interaction prediction losses for $800,000$ steps. The batch size and learning rate are set to $8192$ residues and 33e-44 respectively. Sequences are decoded using the greedy decoding strategy. We incorporate four-level enzyme tags in the EC tree. For example, an enzyme in the category 1.1.1.1 (alcohol dehydrogenase) should have four tags: 1 (oxidoreductases), 1.1 (acting on the CH-OH group of donors), 1.1.1 (with NAD+ or NADP+ as acceptor), and 1.1.1.1 (alcohol dehydrogenase). To further enhance model performance, we first pretrain the model using a masked language modeling (MLM) objective on both sequence and backbone structure, randomly masking 20% of residues and $C_{\alpha}$ coordinates over 800,000 steps. We then follow the standard EnzyGen training procedure. We refer to this variant as EnzyGen-1.5.
@@ 4.2 Experimental Setup
(a) Zero-shot generalization
(b) Fourth-level category clustering
Figure 4: (a) ESP scores of designed enzymes from new fourth-level classes or with new substrates. Dash line denotes median. (b) Fourth-level category embedding clustering.
@@ 4.3 Main Results
Enzyme Family 1.1.1 $1.1$ 1.14.13 1.14.14 1.2.1 2.1.1 2.3.1 2.4.1 2.4.2 2.5.1 2.6.1 2.7.1 2.7.10 2.7.11 2.7.4
@@
EnzyGen excels in designing enzymes that bind their respective substrates with high affinities. Table 2 and 3 highlight the ability of EnzyGen in achieving the highest average ESP score and substrate binding affinity across 323 families. Notably, EnzyGen obtains an average ESP score of $0.65$, surpassing the suggested enzyme-substrate interaction threshold of 0.6 from the ESP evaluator developer. In comparison, PROTSEED achieves 0.59, RFDiffusion+ProteinMPNN achieves 0.53, and ESM2+EGNN achieves 0.61 of average ESP score. EnzyGen’s ESP score is significantly higher than all previous baselines. This is further verified by the binding affinity scores. The enzymes generated by our EnzyGen exhibit stronger binding affinity to their corresponding substrates than all other methods across all 30 third-level testing EC categories. EnzyGen gains an average improvement of 0.87-0.87 in binding affinity over the previous best method (RFDiffusion+ProteinMPNN), which is significant. These results demonstrate that enzymes designed by our EnzyGen are able to bind their corresponding substrates with high affinities. Using the MLM pretraining, the model demonstrates further performance gains. EnzyGen-1.5 achieves superior average enzyme-substrate binding scores across both ESP and docking metrics, although its average pLDDT is marginally lower than that of EnzyGen. These results validate the effectiveness of the pretraining strategy.
5 Analysis: Diving Deep into EnzyGen
(a) Complex of 1KAG-ATP(-4)
(b) Complex of 5L2P-paraoxon
Figure 5: Case Study: (a) Complex of designed 1KAG (2.7.1.71, catalyzing the specific phosphorylation of the 3-hydroxyl group of shikimic acid) and substrate ATP(-4), with pLDDT=90.39, Uniprot blastp recovery rate = 58.5%, (b) Complex of designed 5L2P (3.1.1.2, hydrolyzing various p-nitrophenyl phosphates, aromatic esters and p-nitrophenyl fatty acids) and substrate paraoxon, with pLDDT=89.44, Uniprot blastp recovery rate = 49.4%. Both cases show polar contacts (hydrogen bonds) depicted in purple.
Enzyme Family 1.1.1 $1.1$ 1.14.13 1.14.14 1.2.1 2.1.1 2.3.1 2.4.1 2.4.2 2.5.1 2.6.1 2.7.1 2.7.10 2.7.11 2.7.4
EnzyGen-1.5 2.38 1.04 0.73 1.28 4.08 4.04 5.29 2.98 3.48 3.32 3.22 3.64 3.36 3.35 2.08
Enzyme Family 2.7.7 3.1.1 3.1.3 3.1.4 3.2.2 3.4.19 3.4.21 3.5.1 3.5.2 3.6.1 3.6.4 3.6.5 4.1.1 4.2.1 4.6.1 Avg
EnzyGen-1.5 6.98 3.23 9.72 2.14 7.43 9.18 8.42 3.88 2.03 3.60 9.01 6.92 7.56 1.32 6.20 4.40
Table 6: Self-consistency (RMSD between the folded structure of the designed sequence and the designed structure, $\downarrow$) of EnzyGen-1.5.
(a) 1AZ2, RMSD=0.386Å, EC=1.1.1.21
(b) 1E4V, RMSD=0.619Å, EC=2.7.4.3
(c) 1M9Q, RMSD=0.303Å, EC=1.14.13.39
(d) 4HCU, RMSD=0.560Å, EC=2.7.10.2
Figure 6: Case study of designed backbone structure aligned with ground truth.
@@ 5.6 Designing Novel Enzymes with EnzyGen
EnzyGen is able to design novel enzymes that bind to specific substrates. Figure 5 (a) and (b) showcase two designed enzymes based on PDB 1KAG (catalyzing the specific phosphorylation of the 3-hydroxyl group of shikimic acid) and 5L2P (hydrolyzing various p-nitrophenyl phosphates, aromatic esters and p-nitrophenyl fatty acids) respectively. Both of the two designed enzymes achieve pLDDT approaching or exceeding 90 and substrate binding affinities below -10. Additionally, the Gnina docked complexes show polar contacts (hydrogen bonds) between the enzymes and substrates. This affirms the enzyme-substrate interaction function. We carry out blastp search of designed enzymes in Uniprot, yielding an amino acid identity of 58.5% and 49.4%, respectively, to the most similar enzyme. Together, these results show that our model is capable of designing novel enzymes with high substrate binding affinities. Additional cases are detailed in Appendix D.3, each exhibiting a pLDDT higher than 80, a binding affinity below -10, and a Uniprot blastp identity rate lower than 65%. Enzyme sequences for all analyzed cases are also available in Appendix Table 20.
5.7 How Consistent Are the Designed Sequence and Structure?
To assess the consistency between the designed enzyme sequences and the corresponding backbone structures, we report the RMSD between the folded structure of each designed sequence and the designed structure in Table 6. The results show that 9 out of 30 categories achieve an RMSD below 3, indicating good structural alignment. Additionally, Figure 6 presents four representative cases, each exhibiting an RMSD below 1 between the folded and designed structures, demonstrating near-perfect alignment. These results confirm that EnzyGen-1.5 effectively designs sequences and structures that are highly consistent with one another.
@@ A.1 Proof: NAEL Is SE(3) Equivariant
[…]
@@
We have proved that when $l=0$, $\boldsymbol{h}^{1}_{i}$ is invariant to rotation and translation on $\boldsymbol{x}$. Taking $\boldsymbol{H}^{1}$ as the second layer input and following the above process, we can prove $\boldsymbol{h}^{2}_{i}$ is also invariant. Repeating this process from $l=1$ to L1L-1, we can get the same conclusion.
Combining the above two scenarios together, we have $\boldsymbol{H}^{l+1},R\boldsymbol{x}^{l+1}+\boldsymbol{t}=\mathrm{NAEL}(\boldsymbol{H}^{l},R\boldsymbol{x}^{l}+\boldsymbol{t})$ for $l=0$ to L1L-1. Therefore, our proposed NAEL is $SE(3)$-equivariant.